Two bodies A and B of equal masses moving with velocities (10km/s, 030°) and (15km/s, 090°) impinges and coalesced. Find the initial velocity with which the composite body begins to move immediately after the collision


1
Expert's answer
2021-05-19T20:44:25-0400

Following the conservation of momentum concept;


"m\bold{v_1}+m\bold{v_2}=2m\bold{v}""0.5(10\cos{30},10\sin{30})+0.5(0,15)=(v_x,v_y)""v_x=4.33\frac{m}{s}, v_y=10\frac{m}{s}"

Therefore,



"v=\sqrt{4.33^2+10^2}=10.9\frac{m}{s}""v=\sqrt{4.33^2+10^2}=10.9\frac{m}{s}""\theta=\arctan{\frac{10}{4.33}}=66.6\degree"